64. A museum charges an admission fee of $7.50 per adult and $5.50 per child. On Sunday 300 adults and children visited The museum. A Total of $2000 was collected in admissions. Determine the number of adults and The number of childern that visited The museum on Sunday. Show The steps of your solution.
Answer by macston(5194)

.
C=number of children; A=number of adults
.
A+C=300
A=300-C
.
$7.50(300-C)+$5.50C=$2000
$2250-$7.50C+$5.50C=$2000
-$2.00C=-$250
C=125
ANSWER 1: 125 children visited the museum.
.
A=300-C=300-125=175
ANSWER 2: 175 adults visited the museum.
.
CHECK:
$7.50A+$5.50C=$2000
$7.50(175)+$5.50(125)=$2000
$1312.50+$687.50=$2000
$2000=$2000